# Towards the Worst-case Robustness of Large Language Models

### v1

@@ B.5 On Discrete Diffusion Models
Discrete diffusion models extend traditional diffusion models to the discrete domain, enabling the modeling of language inputs (Meng et al., 2022; Campbell et al., 2022; Lou et al., 2023). Given a vocabulary $\mathcal{V}=\{1,\cdots,|V|\}$, sequence length $N$, a data distribution $p:=p_{0}\in\mathcal{V}^{N}$, the forward process creates a sequence of distributions $p_{t}$ by randomly perturbing each word according to a continuous-time Markov chain:
@@
The forward process has an analytical form due to its simplicity. For the $i$-th word $\bm{x}_{0}^{i}$, $p_{t|0}(\cdot|\bm{x}_{0}^{i})=\exp(\int_{0}^{t}\sigma(s)dsQ)_{\bm{x}_{0}^{i}}$. It also has a well-known reversal given by another diffusion matrix $\overline{Q}_{t}$ (Kelly, 2011). For the $i$-th word, the reversal is:
@@
where $\bm{y}$ is another sentence that differs from $\bm{x}_{t}$ only at $i$-th position, $\frac{p_{t}(\bm{y})}{p_{t}(\bm{x})}$ is referred as the concrete score. Once we train a score network $s_{\theta}(\bm{x}_{t},t)$ to approximate the concrete score, we can sample new instances using Eq. 5 by substituting the unknown score $\frac{p_{t}(\bm{y})}{p_{t}(\bm{x})}$ with the neural network-estimated score $s_{\theta}(\bm{x},t)$ (Meng et al., 2022; Lou et al., 2023). Unlike the forward process, the reverse process lacks an analytical form due to the involvement of a neural network. Consequently, numerical methods such as an Euler solver or a $\tau$-leaping solver are typically employed to approximate the backward Markov chain.
Figure 1: Comparison between DiffPure (Nie et al., 2022) and DiffTextPure using diffusion language models.
@@ B.6 Certification on Different Tasks
Case 1: Image Classification. In image classification, $f$ can be a classifier mapping from the image domain to one interested class in $K-1$ probability simplex. The smoothing distribution $p(\bm{z}|\bm{x})$ can be a Gaussian distribution (Cohen et al., 2019; Chen et al., 2024b), a Uniform distribution (Levine & Feizi, 2021; Lee et al., 2019), Laplacian distribution (Teng et al., 2020), or other types of distributions. If we can certify that $p_{adv}\geq 0.5$ in Definition 5.1, it guarantees that the classifier will consistently produce the correct result for all $\bm{x}_{adv}$ satisfying $\mathcal{D}(\bm{x},\bm{x}_{adv})\leq d$. This is because the probability of the true class remains the highest among all output probabilities.
Case 2: Text Classification. Similarly, for a text classifier $f:\mathcal{V}^{N}\to[0,1]$, that maps from a text to a probability of outputting a target class that we are interested in, the smoothing distribution can be derived from the noisy process of diffusion models, $p_{t|0}(\bm{x}_{\tau}|\bm{x})$, such as randomly replacing or masking words (Lou et al., 2023), as described in Sec. 4. If we can certify that $p_{adv}\geq 0.5$ for the correct class $y$, it ensures that $y$ remains the largest output of $g(\bm{x}_{adv})$, guaranteeing robust classification.
@@
Appendix C Proofs
@@
The optimality of the greedy algorithm in Theorem 5.3 has been extensively proven (Aho & Hopcroft, 1974; Cormen et al., 2022). The proof is typically conducted by contradiction. By sorting the items by their value-to-weight ratio, assume that there exists a better selection than the one obtained by selecting items based on their value-to-weight ratios. Comparing the differing items in these two selections, both must have the same volume, but the selection based on value-to-weight ratio will always have a higher ratio, and thus a higher value. Therefore, in the fractional knapsack problem, it is impossible to find a better approach than selecting items in descending order of their value-to-weight ratio.
Another proof, more closely related to the approach in Cohen et al. (2019), uses the method of Lagrange multipliers. Our goal is to find the minimal solution to a constrained optimization problem:
@@
Further narrowing of the hypothesis set can yield better solutions for this constrained optimization problem, e.g., restricting to binary functions =$\mathcal{F}=\{f:\mathcal{X}\to\{0,1\}\}$ or functions with Lipschitz continuity (Chen et al., 2024a; Delattre et al., 2024).
C.2 Proof of Theorem 5.6
@@
These two combinations are actually the same, as shown by the symmetrization lemma in Theorem C.3. This symmetry provides many favorable properties for the uniform kernel.
@@ C.2.3 Case study: $d=3$
This simple result enables us to intuitively illustrate the greedy algorithm using $p_{adv}-p_{A}$ graph. See Sec. C.4 and Figure 2(a) for detail.
One can also interpret the certified bound for absorbing kernel in another way, similar to Zeng et al. (2023):
@@
C.4 Analytic Solution of Certified Robustness using Absorbing Kernel
(a) $p_{adv}-p_{A}$ plot for the absorbing kernel.
(b) $p_{adv}-p_{A}$ plot for the uniform kernel.
(c) $p_{adv}-p_{A}$ plot for different vocabulary size.
@@ C.4 Analytic Solution of Certified Robustness using Absorbing Kernel
C.5.1 Final Proof of Sufficient Condition Eq. 6
@@
C.5.2 Simple case study: $|\mathcal{V}|\to\infty$
@@
When $|\mathcal{V}|\to\infty$, the value-to-weight ratio $\frac{\alpha^{i}\bar{\beta}^{d-i}}{\alpha^{j}\bar{\beta}^{d-j}}=\frac{\alpha^{i-j}}{\bar{\beta}^{i-j}}=\left(\frac{\alpha}{\bar{\beta}}\right)^{i-j}$ only have three possible values: 0 when $i>j$, 1 when $i=j$, $\infty$ when $i<j$. Since for all $p_{A}\leq 1-\beta^{d}$, we have $i>j$, thus $p_{adv}=0$ for all $p_{A}\leq 1-\beta^{d}$. By symmetrization lemma (Theorem C.3), $i=j$ must hold for all $p_{A}\geq 1-\beta^{d}$. Therefore, the $p_{adv}-p_{A}$ graph of the uniform kernel and absorbing kernel is exactly the same. This means Figure 2(b) gradually goes to Figure 2(a) when $|\mathcal{V}|\to\infty$.
C.5.3 Simple case study: $d$=1,2,3
@@
C.6 Symmetrization Lemma
@@
Through the symmetrization lemma Theorem C.3, we have the following corollary, which will be used in Sec. C.7.
Corollary C.4.
@@
C.7 Relationship Between $|\mathcal{V}|$ and Certified Radius
@@
Using the symmetrization lemma (Theorem C.3), we only need to prove the case where the trading rate $\left(\frac{\alpha}{\bar{\beta}}\right)^{j-i}\leq 1$, i.e., $j\geq i$. In this proof, unless stated otherwise, we assume $j\geq i$.
@@
From Lemma C.2, we have:
@@
C.8 0-1 Knapsack
1:  Let $n$ be the number of items
2:  Initialize DP table $dp[i][w]=-\infty$ for all $i\leq n$ and $w\leq p_{A}$
3:  Set $dp[0][w]=0$ for all $w$, as no items result in zero value
4:  for each item $\bm{z}^{(i)}$ from $1$ to $n$ do
5:     for each possible weight $w\leq p_{A}$ do
6:        Update DP table:
7:     end for
8:  end for
9:  Let $V_{\text{max}}=-\max(dp[n][w])$ for $w=0,1,\dots,W_{\max}$Maximum value obtainable with capacity $W_{\max}$}
10:  Return: $\mathbb{I}\{V_{\text{max}}\geq\tau\}$Return 1 if value $V_{\text{max}}$ is greater than or equal to threshold $\tau$, else return 0}
Appendix D Tricks
D.1 Parameterizing $t$ as $1-\bar{\beta}$
@@
This technique has been extensively discussed in Karras et al. (2022) and Chen et al. (2024b). For clarity, our code also adopts this parameterization.
D.2 LLMs as Detectors
In contrast, prompting LLMs makes debugging significantly easier. For example, we can directly ask the LLM, “Why do you think this sentence is harmful?” and gain insights into its reasoning. This makes the process of debugging and controlling false positive rates much more intuitive and transparent.
We do not adopt Llama-3 Guard (Dubey et al., 2024) in our approach because it exhibits a higher false positive rate compared to our method, primarily due to its non-conservative prompt design.
Appendix E More Experiments Details
E.1 Our (Nearly) 0 False Positive Prompt
@@
DiffTextPure (Uniform) 98% 92% 90% 94% 16% 92%
DiffTextPure (Absorb) 98% 86% 92% 94% 30% 86%
@@
DiffTextPure (Uniform) 100% 98% 100% 100% 100% 100%
DiffTextPure (Absorb) 100% 100% 100% 100% 100% 100%
@@
DiffTextPure (Uniform) 96% 100% 100% 94% 73% 100%
DiffTextPure (Absorb) 96% 100% 100% 98% 69% 100%
E.2 Experimental Detail of Empirical Evaluations
Dataset: Following prior works, we use the AdvBench dataset (Zou et al., 2023), which consists of approximately 500 harmful strings and behaviors. Due to limited computational resources, we follow Jia et al. (2024) and use their harmful behaviors subset, which contains 50 behaviors randomly sampled from AdvBench.
Baselines: We compare our defense against four state-of-the-art baselines—PPL (Alon & Kamfonas, 2023), ICD (Wei et al., 2023b), Self-reminder (Wu et al., 2023), and PAT (Mo et al., 2024) across four types of jailbreak attacks: GCG (Zou et al., 2023), MAC (Zhang & Wei, 2025), I-GCG (Jia et al., 2024), AutoDAN (Liu et al., 2023), ICA (Wei et al., 2023b) and our I2-GCG (see Sec. 3).
Models: Our experiments span four open-source models, including Vicuna-7B (Zheng et al., 2024a), Llama-2-7B-Chat (Touvron et al., 2023), and Llama-3-8B-Instruct (Dubey et al., 2024).
Hyper-parameters: The experimental settings for baseline attacks and defenses follow their original papers, except for two adjustments: we use a 5-shot setting for ICA and optimize for 100 steps in AutoDAN, due to memory constraints. For hyper-parameters in DiffTextPure, we adopt $\beta=0.25$. We use the diffusion language model (Lou et al., 2023) as the purifier.
E.3 Black-box Evaluation
Black-box evaluations represent practical settings where attackers have only limited access to the model. In this section, we follow previous work (Wei et al., 2023b; Mo et al., 2024; Wu et al., 2023) and conduct experiments in which the attackers know only the base model but are unaware of the defense.
@@ E.3 Black-box Evaluation
E.4 Certified Robustness Settings
Following previous work (Cohen et al., 2019; Salman et al., 2019; Carlini et al., 2023b; Xiao et al., 2023; Chen et al., 2024a), we use sample size $1,000,000$, type one error $0.01$. In main experiments, we use $\beta=0.1$ for certification against $\ell_{0}$ attacks, and $\beta=0.25$ for certification against the suffix attacks. We use the diffusion language models (Lou et al., 2023) as the purifier in the main experiments and also compare with the GPT-4o purifier in Sec. E.6.

### v2

@@ B.5 On Discrete Diffusion Models
Discrete diffusion models extend traditional diffusion models to the discrete domain, enabling the modeling of language inputs [54, 118, 20]. Given a vocabulary $\mathcal{V}=\{1,\cdots,|V|\}$, sequence length $N$, a data distribution $p:=p_{0}\in\mathcal{V}^{N}$, the forward process creates a sequence of distributions $p_{t}$ by randomly perturbing each word according to a continuous-time Markov chain:
@@
The forward process has an analytical form due to its simplicity. For the $i$-th word $\bm{x}_{0}^{i}$, $p_{t|0}(\cdot|\bm{x}_{0}^{i})=\exp(\int_{0}^{t}\sigma(s)dsQ)_{\bm{x}_{0}^{i}}$. It also has a well-known reversal given by another diffusion matrix $\overline{Q}_{t}$ [119]. For the $i$-th word, the reversal is:
@@
where $\bm{y}$ is another sentence that differs from $\bm{x}_{t}$ only at $i$-th position, $\frac{p_{t}(\bm{y})}{p_{t}(\bm{x})}$ is referred as the concrete score. Once we train a score network $s_{\theta}(\bm{x}_{t},t)$ to approximate the concrete score, we can sample new instances using Eq. 9 by substituting the unknown score $\frac{p_{t}(\bm{y})}{p_{t}(\bm{x})}$ with the neural network-estimated score $s_{\theta}(\bm{x},t)$ [54, 20]. Unlike the forward process, the reverse process lacks an analytical form due to the involvement of a neural network. Consequently, numerical methods such as an Euler solver or a $\tau$-leaping solver are typically employed to approximate the backward Markov chain.
@@ B.6 Certification on Different Tasks
Case 1: Image Classification. In image classification, $f$ can be a classifier mapping from the image domain to one interested class in $K-1$ probability simplex. The smoothing distribution $p(\bm{z}|\bm{x})$ can be a Gaussian distribution [23, 113], a Uniform distribution [120, 117], Laplacian distribution [27], or other types of distributions. If we can certify that $p_{adv}\geq 0.5$ in Definition 4.1, it guarantees that the classifier will consistently produce the correct result for all $\bm{x}_{adv}$ satisfying $\mathcal{D}(\bm{x},\bm{x}_{adv})\leq d$. This is because the probability of the true class remains the highest among all output probabilities.
Case 2: Text Classification. Similarly, for a text classifier $f:\mathcal{V}^{N}\to[0,1]$, that maps from a text to a probability of outputting a target class that we are interested in, the smoothing distribution can be derived from the noisy process of diffusion models, $p_{t|0}(\bm{x}_{\tau}|\bm{x})$, such as randomly replacing or masking words [20], as described in Sec. F.2. If we can certify that $p_{adv}\geq 0.5$ for the correct class $y$, it ensures that $y$ remains the largest output of $g(\bm{x}_{adv})$, guaranteeing robust classification.
@@
Appendix C Proofs for Knapsack Solvers
@@
The optimality of the greedy algorithm in Theorem 4.3 has been extensively proven [50, 51]. The proof is typically conducted by contradiction. By sorting the items by their value-to-weight ratio, assume that there exists a better selection than the one obtained by selecting items based on their value-to-weight ratios. Comparing the differing items in these two selections, both must have the same volume, but the selection based on value-to-weight ratio will always have a higher ratio, and thus a higher value. Therefore, in the fractional knapsack problem, it is impossible to find a better approach than selecting items in descending order of their value-to-weight ratio.
Another proof, more closely related to the approach in [23], uses the method of Lagrange multipliers. Our goal is to find the minimal solution to a constrained optimization problem:
@@
Further narrowing of the hypothesis set can yield better solutions for this constrained optimization problem, e.g., restricting to binary functions =$\mathcal{F}=\{f:\mathcal{X}\to\{0,1\}\}$ or functions with Lipschitz continuity [52, 53].
C.2 0-1 Knapsack
1:  Let $n$ be the number of items
2:  Initialize DP table $dp[i][w]=-\infty$ for all $1\leq i\leq n$ and $w\leq p_{A}$, set $dp[i][0]=0$ for all $i\leq n$
3:  for each item $\bm{z}^{(i)}$ from $1$ to $n$ do
4:     for each possible weight $w\leq p_{A}$ do
5:        Update DP table:
6:     end for
7:  end for
8:  Let $V_{\text{max}}=-dp[n][w]$
9:  Return: $\mathbb{I}\{V_{\text{max}}\geq\tau\}$Return 1 if value $V_{\text{max}}$ is greater than or equal to threshold $\tau$, else return 0}
Appendix D Proofs For Value-to-weight Ratio and Volume for Specific Kernels
D.1 Proof of Theorem 5.4
@@ Appendix D Proofs For Value-to-weight Ratio and Volume for Specific Kernels
These two combinations are actually the same, as shown by the symmetrization lemma in Theorem E.1. This symmetry provides many favorable properties for the uniform kernel.
@@ D.1.3 Case study: $d=3$
This simple result enables us to intuitively illustrate the greedy algorithm using $p_{adv}-p_{A}$ graph. See Sec. D.3 and Figure 1(a) for detail.
One can also interpret the certified bound for absorbing kernel in another way, similar to [29]:
@@
D.3 Analytic Solution of Certified Robustness using Absorbing Kernel
@@
D.4.1 Final Proof of Sufficient Condition Eq. 10
@@
D.4.2 Simple case study: $|\mathcal{V}|\to\infty$
@@
When $|\mathcal{V}|\to\infty$, the value-to-weight ratio $\frac{\alpha^{i}\bar{\beta}^{d-i}}{\alpha^{j}\bar{\beta}^{d-j}}=\frac{\alpha^{i-j}}{\bar{\beta}^{i-j}}=\left(\frac{\alpha}{\bar{\beta}}\right)^{i-j}$ only have three possible values: 0 when $i>j$, 1 when $i=j$, $\infty$ when $i<j$. Since for all $p_{A}\leq 1-\beta^{d}$, we have $i>j$, thus $p_{adv}=0$ for all $p_{A}\leq 1-\beta^{d}$. By symmetrization lemma (Theorem E.1), $i=j$ must hold for all $p_{A}\geq 1-\beta^{d}$. Therefore, the $p_{adv}-p_{A}$ graph of the uniform kernel and absorbing kernel is exactly the same. This means Figure 1(b) gradually goes to Figure 1(a) when $|\mathcal{V}|\to\infty$.
D.4.3 Simple case study: $d$=1,2,3
@@
Appendix E Reduction Lemma and Symmetrization Lemma
E.1 Reduction Lemma
For value-to-weight ratio:
For its volume:
E.2 Symmetrization Lemma
@@
Through the symmetrization lemma Theorem E.1, we have the following corollary, which will be used in Sec. E.3.
Corollary E.2.
@@
E.3 Relationship Between $|\mathcal{V}|$ and Certified Radius
@@
Using the symmetrization lemma (Theorem E.1), we only need to prove the case where the trading rate $\left(\frac{\alpha}{\bar{\beta}}\right)^{j-i}\leq 1$, i.e., $j\geq i$. In this proof, unless stated otherwise, we assume $j\geq i$.
@@
From Lemma D.1, we have:
@@
Appendix F Tricks
This section presents some “implementation tricks” of previous defenses evaluated in this paper.
F.1 LLMs as Detectors
In contrast, prompting LLMs makes debugging significantly easier. For example, we can directly ask the LLM, “Why do you think this sentence is harmful?" and gain insights into its reasoning. This makes the process of debugging and controlling false positive rates much more intuitive and transparent.
We do not adopt Llama-3 Guard [3] in our approach because it exhibits a higher false positive rate compared to our method, primarily due to its non-conservative prompt design.
F.2 DiffTextPure: Diffuse Text and Purify
To construct a smooth function $g(\bm{x})=\mathbb{E}_{p(\bm{z}|\bm{x})}[f(\bm{z})]$ that possesses theoretical guarantees, we first need to apply a forward process to $\bm{x}$, generating a noised sample $\bm{z}\sim p(\bm{z}|\bm{x})$, e.g., Absorbing kernel, which replaces each token with a mask token with probability $\beta$; Uniform kernel, which replaces each token with another token from the vocabulary uniformly at random with probability $\beta$.
[…]
Algorithm 3 DiffTextPure
  Sample $\bm{z}\sim p(\bm{z}|\bm{x}_{adv})$, or for specific case:
  if $p(\bm{z}|\bm{x})$ is Absorb then
     Mask tokens with probability $\beta$.
  else if $p(\bm{z}|\bm{x})$ is Uniform then
     Replace tokens uniformly with probability $\beta$.
  end if
  Purify $\bm{z}$ using a LLM or a diffusion model to obtain $\bm{x}_{0}$.
  Return: The output of sub-sequent function $\hat{f}(\bm{x}_{0})$.
F.2.1 Understanding DiffTextPure
Lemma F.1 (DiffTextPure increases the likelihood).
We do two case studies in the following:
Case I: destroying adversarial strings. Since the forward process randomly perturbs each word, DiffTextPure has a high probability of disrupting the adversarial nature of the adversarial suffix. As demonstrated in the example below, frequently occurring words in adversarial suffixes, such as “Wikipedia"—which often encourages language models to respond to harmful queries similarly to a Wikipedia article [10]—are replaced with non-harmful symbols, thus neutralizing their adversarial impact. Although the user’s request may also be perturbed, the reverse denoising process allows the denoiser $D$ to recover this part due to its high likelihood, while the adversarial suffix is nearly impossible to be restored due to the extremely low likelihood of encountering such garbled strings.
Figure 2: Comparison between DiffPure [106] and DiffTextPure using diffusion language models.
{mdframed}
[backgroundcolor=gray!20, linewidth=1pt, roundcorner=10pt]
[…]
Assistant: Sorry, I can’t assist with that.
Case II: purifying into benign requests. Since adversarial requests are typically absent from the training set of denoiser $D$, the reverse process is more likely to recover the noisy text as a benign request rather than a harmful one. As demonstrated in the example below, DiffTextPure changes the phrase “hack into" to “log into," effectively defending against harmful requests.
{mdframed}
[backgroundcolor=gray!20, linewidth=1pt, roundcorner=10pt]
Assistant: Sure, here is how to log into a corporate network.
F.3 Parameterizing $t$ as $1-\bar{\beta}$
@@
This technique has been extensively discussed in Karras et al. [121] and Chen et al. [113]. For clarity, our code also adopts this parameterization.
Appendix G More Experiments Details
G.1 Our (Nearly) 0 False Positive Prompt
@@
G.2 Experimental Detail of Empirical Evaluations
Dataset: Following prior works, we use the AdvBench dataset [10], which consists of approximately 500 harmful strings and behaviors. Due to limited computational resources, we follow Jia et al. [17] and use their harmful behaviors subset, which contains 50 behaviors randomly sampled from AdvBench.
Baselines: We compare our defense against four state-of-the-art baselines—PPL [40], ICD [36], Self-reminder [42], and PAT [41] across four types of jailbreak attacks: GCG [10], MAC [72], I-GCG [17], AutoDAN [39], ICA [36] and our I2-GCG (see Sec. 3).
Models: Our experiments span four open-source models, including Vicuna-7B [58], Llama-2-7B-Chat [122], and Llama-3-8B-Instruct [3].
Hyper-parameters: The experimental settings for baseline attacks and defenses follow their original papers, except for two adjustments: we use a 5-shot setting for ICA and optimize for 100 steps in AutoDAN, due to memory constraints. For hyper-parameters in DiffTextPure, we adopt $\beta=0.25$. We use the diffusion language model [20] as the purifier.
@@ G.2 Experimental Detail of Empirical Evaluations
Uniform 98% 92% 90% 94% 16% 92%
Absorb 98% 86% 92% 94% 30% 86%
@@
Uniform 100% 98% 100% 100% 100% 100%
Absorb 100% 100% 100% 100% 100% 100%
@@
Uniform 96% 100% 100% 94% 73% 100%
Absorb 96% 100% 100% 98% 69% 100%
G.3 Black-box Evaluation
Black-box evaluations represent practical settings where attackers have only limited access to the model. In this section, we follow previous work [36, 41, 42] and conduct experiments in which the attackers know only the base model but are unaware of the defense.
@@ G.3 Black-box Evaluation
G.4 Certified Robustness Settings
Following previous work [23, 45, 111, 109, 52], we use sample size $1,000,000$, type one error $0.01$. In main experiments, we use $\beta=0.1$ for certification against $\ell_{0}$ attacks, and $\beta=0.25$ for certification against the suffix attacks. We use the diffusion language models [20] as the purifier in the main experiments and also compare with the GPT-4o purifier in Sec. G.6.